# SCHEME: Scalable Channer Mixer for Vision Transformers

### v1

@@ Abstract
Vision Transformers have received significant attention due to their impressive performance in many vision tasks. While the token mixer or attention block has been studied in great detail, the channel mixer or feature mixing block (FFN or MLP) has not been explored in depth albeit it accounts for a bulk of the parameters and computation in a model. In this work, we study whether sparse feature mixing can replace the dense connections and confirm this with a block diagonal MLP structure that improves the accuracy by supporting larger expansion ratios. To improve the feature clusters formed by this structure and thereby further improve the accuracy, a lightweight, parameter-free, channel covariance attention (CCA) mechanism is introduced as a parallel branch during training. This design of CCA enables gradual feature mixing across channel groups during training whose contribution decays to zero as the training progresses to convergence. This allows the CCA block to be discarded during inference, thus enabling enhanced performance with no additional computational cost. The resulting Scalable CHannEl MixEr (SCHEME) can be plugged into any ViT architecture to obtain a gamut of models with different trade-offs between complexity and performance by controlling the block diagonal structure size in the MLP. This is shown by the introduction of a new family of SCHEMEformer models. Experiments on image classification, object detection, and semantic segmentation, with different ViT backbones, consistently demonstrate substantial accuracy gains over existing designs, especially under lower FLOPs regimes. For example, the SCHEMEformer establishes a new SOTA of 79.7% accuracy for ViTs using pure attention mixers on ImageNet-1K at 1.77G FLOPs.
@@ 1 Introduction
Figure 1: Left: Accuray-FLOPs tradeoff for various SCHEME models (denoted in blue) based on popular ViT architectures. Right: Accuracy-FLOPs trade-off of the proposed SCHEMEformer family derived from the Metaformer-PPAA-S12 model [57] with higher expansion ratios in the MLP blocks. The SCHEMEFormer family outperforms the baseline for the same FLOPs, with larger gains for smaller models.
[…]
@@
Vision Transformers (ViTs) [15, 33, 51, 32] have become ubiquitous in computer vision in recent years. Under this architecture, an image is decomposed into a set of patches which are considered as tokens and fed to a transformer model [49] of two main components: a spatial attention module, which reweighs each token according to its similarity to the other tokens extracted from the image, enabling information fusion across large spatial distances, and a channel mixer that combines the feature channels extracted from all patches using a multi-layer perceptron (MLP or FFN). A bottleneck of this model is the quadratic complexity of the attention mechanism on the number of patches. Numerous ViT variants have been proposed in recent years to improve complexity and/or accuracy. These can be grouped into 1) improved attention mechanisms, and 2) hybrid architectures that replace attention or combine it with convolutions.
Much less research has been devoted to the channel mixer. Most models simply adopt the MLP block of [49], consisting of two layers, where channels are first expanded by a specified expansion ratio and then compressed to the original dimension. This is somewhat surprising since the mixer is critical for good transformer performance. For example, Dong et. al.  [14] showed that pure attention, without MLPs or residual connections, collapses doubly exponentially to a rank one matrix. Yu et. al. [57] also showed that without residual connections or MLP, the model cannot converge. DaViT [12] showed that replacing MLPs with more attention blocks (both spatial and channel attention) of equivalent computational complexity lowers the transformer accuracy. These observations show that the channel mixer is an indispensable component of ViTs.
In this work, we investigate the channel mixer module, both to quantify its contribution to ViT performance and to improve the complexity vs. accuracy trade-off of the transformer model. We start by studying the impact of the channel expansion performed in the mixer, by varying its expansion ratio. Transformer performance is shown to increase until it saturates for an expansion ratio beyond 8. However, because mixer (MLP) complexity is linear on the dimension of the intermediate representation, naive channel scaling with expansion ratios larger than 4 causes an explosion of parameters and computation, leading to models of large complexity and prone to overfit.
To achieve a better trade-off between dimensionality and computation, we seek to leverage a hardware efficient mixer with block diagonal structure [1, 10]. This consists of breaking the input and output feature vectors of a network layer into disjoint groups and performing matrix multiplications only within each group, as illustrated in Figure 2. We denote the resulting MLP block as Block Diagonal MLP (BD-MLP). Despite the lack of feature mixing across groups, we find that a transformer model equipped with the BD-MLP and a larger expansion ratio (to match both the parameter count and computation) achieves comparable or slightly higher accuracy than a baseline with the dense MLP. This suggests that the lower accuracy of the sparse block diagonal operations is offset by the gains of larger expansion ratios. Further analyzing the features learned by the different groups of the BD-MLP, we observe that they form feature clusters that are similar in terms of their discriminative ability to identify the target classes.
To learn better feature clusters, we seek a mechanism capable of restoring inter-group feature communication only during training without increasing parameters and computational overhead. For this, we propose to use the channel correlations among features. This is done by introducing a channel attention branch that computes the channel covariance matrix of the input features of the BD-MLP and is used to re-weigh its output features, as illustrated in Figure 2. This attention mechanism is denoted as the channel covariance attention (CCA) block. The reweighted features are then fused with the BD-MLP output by means of a weighted residual addition with learned weights $(\alpha,1-\alpha)$. The overall design, combining the sparse block diagonal structure of the BD-MLP, and the parameter-free CCA attention module, is extremely efficient in terms of parameters. Experiments also show that the fusion weight learned in the CCA branch, denoted as $1-\alpha$ in Fig. 2, gradually decays to zero during training. This happens consistently across all layers of the model and across model architectures. While the CCA is important for the formation of good feature clusters during training, it can be removed at inference without any loss, as illustrated in Figure 2. As a result, the model accuracy improves over simply using the BD-MLP mixer, but the computational complexity of model inference does not. This leads to an extremely efficient inference setup, both in terms of parameters and FLOPs. We empirically verify the hypothesis through ablations that study the evolution of attention branch weights over training and the class separability of the learned feature groups.
Since the combination of BD-MLP and CCA enables the design of models with larger expansion ratios, we denote the resulting mixer as the Scalable CHannEl MixEr (SCHEME). SCHEME is a generic channel mixer that can be plugged into existing ViT variants to realize larger expansion ratios and obtain effective scaled-down or scaled-up model versions. We replace the channel mixer of the MetaFormer-PPAA-S12 [57] model with SCHEME to obtain a new family of SCHEMEformer models with improved accuracy/complexity trade-off. This is shown in the right of Figure 1. For the same parameter count and FLOPs, SCHEME obtains better accuracy than the baseline by using larger MLP expansion ratios. The gains are particularly larger for the smaller models. For example, the SCHEMEformer family achieves SOTA results of 79.74% accuracy at 1.77 G FLOPs, for ViTs using self-attention mixers. The left of Figure 1 shows that SCHEME models based on popular ViT architectures have a better Accuracy-FLOPs trade-off than the baseline models. Most of the SCHEMEformer models lie on the Pareto frontier of accuracy vs complexity, showing that SCHEME allows fine control over the two variables, while guaranteeing SOTA performance. These properties are shown to extend to object detection and semantic segmentation tasks, as well as to different architectures, such as T2T-ViT [60], CoAtNet [9], Swin Transformer [32], CSWin [13] and DaViT [12]. In summary, our contributions are as follows,
@@
  • We study the channel mixer of MLPs in ViT and show that the dense feature mixing can be replaced by sparse feature mixing with higher internal feature dimensionality for improved accuracy.
@@
  • We propose the SCHEME module, combining 1) a BD-MLP to enable internal feature representations of larger dimensionality than previous MLP blocks, and 2) CCA to enable the learning of these representations without cost at inference.
@@
  • Experiments on image classification, object detection, and semantic segmentation consistently show gains in accuracy for fixed computation and model size.
@@ 2 Related Work
Vision Transformers: Vision transformers advanced the state-of-the-art in several vision tasks since [33, 15] successfully applied the transformer-based self-attention NLP model of [49] to image generation and classification tasks. These models rely on a spatial attention mechanism, based on the matrix of dot-products between features extracted from image patches. This has quadratic complexity in the number of patches and is quite intensive. Most follow up work [19, 40, 59, 45, 41, 54, 43, 60, 62, 50] improved the spatial attention mechanism of ViT. DeiT [43] and subsequent works [46, 47] introduced a distillation token to distill information, typically from a CNN teacher, without large amounts of data or compute. PvT [51] proposed a progressive shrinking pyramid architecture with spatial-reduction attention that scales ViTs for dense prediction tasks beyond image classification. Swin transformers [32] introduced a hierarchical shifted window attention mechanism, which reduces the complexity to linear with respect to the number of windows. HaloNet [50] proposed two extensions for local ViTs [34] with blocked local attention and relaxed translational equivariance for scaling ViTs. A more extensive review of ViTs is given in [31].
@@
Recently, several works have shown that spatial attention is not the critical ViT feature. Some works improved performance by relying on hybrid architectures, which augment or replace ViT layers with convolutions [53, 48]. Efficient transformer designs, for edge devices, frequently sacrifice spatial attention to achieve better trade-offs between FLOPs and accuracy [5, 27]. Other works have questioned the need for spatial attention altogether. Metaformer [57] argued that the fundamental trait of ViT is the mixing of information across patches, showing that competitive results can be obtained by simply replacing attention with pooling or identity operations. Similarly, [30] showed that spatial attention is not critical for vision transformers by proposing a spatial-gating MLP of comparable performance to ViT [15]. DaViT [12] showed that it is helpful by simply reusing the spatial attention design for channel attention, building a cascade of alternating spatial and channel attention blocks. While this improves performance, it increases the complexity of the transformer block, resulting in a model with many parameters and potentially redundant channel mixer operations. Despite all this work on ViT architectures, little emphasis has been devoted to the other main component of the transformer: the channel mixer module (MLP) that follows attention. This is surprising because the mixer dominates both the parameter count and complexity (FLOPs) of the standard transformer block. [26, 59] modify the MLP block to mimic the inverted residual block of the MobileNetV2 [36], by adding a depthwise convolution. This improved performance but increases parameter and computation costs. Switch Transformer [17] replaces the FFN with sparse mixture of experts [39] that dynamically routes the input tokens. This design allows scaling models to large sizes using higher number of experts, however its not effective for ViTs. In this work, we instead propose an efficient and generic channel mixer module that improves both the parameter and computational efficiency of the transformer and allows for flexible scaling of ViTs.
@@ 3 The SCHEME module
Figure 2 depicts the proposed SCHEME module for feature mixing in ViTs. As shown in the figure, the standard channel mixer consists of two MLP layers, which expand the dimensionality of the input features and then reduce it to the original size. Let $\mathbf{x}\in\mathbb{R}^{d\times N}$ be the matrix containing the $N$ $d$-dimensional input feature vectors extracted from $N$ image patches. The mixer computes an intermediate representation $\mathbf{z}\in\mathbb{R}^{Ed\times N}$ and an output representation $\mathbf{y}\in\mathbb{R}^{d\times N}$ according to
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{Ed\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times Ed},\mathbf{b}_{1}\in\mathbb{R}^{d},\mathbf{b}_{2}\in\mathbb{R}^{Ed},$ ${\bf 1}_{N}$ is the $N$-dimensional vector containing ones as all entries, $\sigma(.)$ is the activation function (typically GELU), and $E$ is an expansion factor, typically 4.
Table 1 details the impact of the mixer on overall transformer performance by evaluating the role of the expansion factor $E$ on the performance of MetaFormer-PPAA-S12 [57] (Pooling, Pooling, Attention, Attention) architecture on ImageNet-1K. Classification accuracy increases from 76.0 to $81.8\%$, as $E$ ranges from $1$ to $6$, decreasing for $E=8$, which suggests overfitting. The table also shows that these gains are not trivial. The MetaFormer-S18, which has no expansion ($E=1$) but more transformer layers and complexity comparable to that of the MetaFormer-S12 ($E=4$), has a significantly lower accuracy of 78.3%. In summary, for a given computation budget, it is beneficial to trade off transformer depth for dimensionality expansion in the channel mixer. This shows that this expansion is a critical component of the transformer architecture. On the other hand, naively scaling $E$ beyond 6 severely increases parameters and computation, leading to models that over-fit and are impractical for many applications.
@@ 3.1 Scalable Channel Mixer (SCHEME)
To efficiently realize the gains of higher intermediate MLP feature dimensionality, we propose the SCHEME module, which consists of a BD-MLP and CCA block where the CCA branch is only used during training. The following sections describe these in detail.
Block Diagonal MLP (BD-MLP): Block diagonal matrices have been previously proposed to efficiently approximate dense matrices [1, 10]. In CNNs, group channel operations are frequently used to design lightweight mobile models with improved accuracy-computation trade-off [25, 6, 36]. This consists of splitting the feature vectors of (1)-(2) into disjoint groups, e.g. $\mathbf{x}$ into a set of $g_{1}$ disjoint features $\{\mathbf{x}_{k}\}_{k=1}^{g_{1}}$ where $\mathbf{x}_{k}\in\mathbb{R}^{d/g_{1}\times N}$, and $\mathbf{y}$ into a set $\{\mathbf{y}_{k}\}_{k=1}^{g_{2}}$ where $\mathbf{y}_{k}\in\mathbb{R}^{Ed/g_{2}\times N}$. As illustrated in Figure 2, the MLPs of (1)-(2) are then implemented independently for each group, according to
@@
$\mathbf{y}_{k}=\mathbf{W}_{2.k}\mathbf{z}_{k}+\mathbf{b}_{2,k}{\bf 1}^{T}_{N}$ (4)
where $\mathbf{W}_{1}\in\mathbb{R}^{Ed/g_{1}\times d/g_{1}},\mathbf{W}_{2}\in\mathbb{R}^{d/g_{2}\times Ed/g_{2}},\mathbf{b}_{1}\in\mathbb{R}^{d/g_{1}},\mathbf{b}_{2}\in\mathbb{R}^{Ed/g_{2}}$ and $\mathbf{z}$ is decomposed into a set $\{\mathbf{z}_{k}\}_{k=1}^{G}$ where $\mathbf{z}_{k}\in\mathbb{R}^{d/G\times N}$, with $G=g_{1}$ in (3) and $G=g_{2}$ in (4). Since the complexity of (3) is $g_{1}^{2}$ times smaller than that of (1) and there are $g_{1}$ groups, the complexity of the first MLP is $1/g_{1}$ times that of standard MLP. Similarly, the complexity of the second MLP is $1/g_{2}$ times that of the standard MLP. Hence, a transformer equipped with the BD-MLP and expansion factor $\frac{2g_{1}g_{2}}{g_{1}+g_{2}}E$ has identical complexity to a standard transformer of factor $E$. For example, when $g_{1}=g_{2}=g$ this allows growing the expansion factor by a factor of $g$ without computational increase.
Expansion FLOPs Top-1 Acc
Model ratio ($E$) (G) (%)
[…]
Metaformer-PPAA-S12 [57] 4 2.55 81.0 (+0.0)
[…]
Table 1: Effect of the MLP block expansion ratio ($E$) on ImageNet-1K validation accuracy for the MetaFormer-S12 [57] architecture. *{*}* denotes results obtained with the code made publicly available by the authors.
@@
where the softmax operation is applied across the matrix rows and $\tau$ is a smoothing factor. The output of the channel mixer block is obtained by the weighted summation of the BD-MLP and CCA branches, according to
$\begin{split}\textbf{y}_{out}=\alpha\textbf{y}+(1-\alpha)CCA(\textbf{x})\end{split}$ (6)
@@
CCA as a Regularizer: The design choice of using a parameter free attention branch and a static weight $\alpha$ allows the model to form better feature clusters during training and gradually decay the contribution from CCA branch once the feature clusters are formed. This can be seen in Figure 3(a), which plots the value of the learned mixing weight $1-\alpha$ as a function of training epochs on ImageNet-1K, for all transformer layers. These plots are typical of the behavior we observed with all transformer backbones and architectures we considered. Clearly, $1-\alpha$ starts with high to intermediate values, indicating that information flows through both branches of the mixer, but decays to $1-\alpha\approx 0$ as training converges. Hence, there is no degradation if the CCA branch is removed during inference. This eliminates a substantial amount of computation during inference, leading to the the training and inference setup of Figure 2, where CCA is not used at inference.
We explain this behavior by conjecturing that the downside of the computational efficiency of the BD-MLP is a more difficult learning problem, due to the independent processing of channel groups. This creates symmetries in the cost function, e.g. the order of the feature groups is not important, and requires a feature clustering operation that is likely to produce more local minima. The CCA branch helps to smooth out this cost function during training, while the feature groups are not established, by allowing inter-group communication. However, once the right feature groupings are found, CCA is no longer needed, and a simple BD-MLP mixer has no loss of performance over the standard MLP. Note that the operation of (5) is basically a projection of $\mathbf{x}$ into canonical subspaces of features that are correlated in the input image. This is likely to be informative to guide the group formation, but less useful when the features are already clustered. While this hypothesis is not trivial to test, since (5) varies from example to example, we confirmed that using CCA during training enhances class separability, which likely reduces overfitting for large expansion ratios. See the results section for more details.

### v4

# SCHEME: Scalable Channel Mixer for Vision Transformers

@@ Abstract
Vision Transformers have achieved impressive performance in many vision tasks. While the token mixer or attention block has been studied in great detail, much less research has been devoted to the channel mixer or feature mixing block (FFN or MLP), which accounts for a significant portion of the model parameters and computation. In this work, we show that the dense MLP connections can be replaced with a sparse block diagonal MLP structure that supports larger expansion ratios by splitting MLP features into groups. To improve the feature clusters formed by this structure we propose the use of a lightweight, parameter-free, channel covariance attention (CCA) mechanism as a parallel branch during training. This enables gradual feature mixing across channel groups during training whose contribution decays to zero as the training progresses to convergence. As a result, the CCA block can be discarded during inference, enabling enhanced performance at no additional computational cost. The resulting Scalable CHannEl MixEr (SCHEME) can be plugged into any ViT architecture to obtain a gamut of models with different trade-offs between complexity and performance by controlling the block diagonal MLP structure. This is shown by the introduction of a new family of SCHEMEformer models. Experiments on image classification, object detection, and semantic segmentation, with 12 different ViT backbones, consistently demonstrate substantial accuracy/latency gains (upto 1.5% /20%) over existing designs, especially for lower complexity regimes. The SCHEMEformer family is shown to establish new Pareto frontiers for accuracy vs FLOPS, accuracy vs model size, and accuracy vs throughput, especially for fast transformers of small size.
@@ 1 Introduction
Refer to caption Refer to caption Refer to caption
(a) Accuracy-FLOPs (b) Accuracy-Parameters (c) Accuracy-Throughput
Figure 1: Comparison of the proposed SCHEMEformer family, derived from the Metaformer-PPAA-S12 model [55] with higher expansion ratios in the MLP blocks, and many SOTA transformers from the literature. The SCHEMEFormer family establishes a new Pareto frontier (optimal trade-off) for a) accuracy vs. FLOPs, b) accuracy vs model size, and c) accuracy vs, throughput. SCHEMEformer models are particularly effective for the design of fast transformers (throughput between 75 and 150 images/s) with small model size. See Appendix Figure 9 for zoomed version.
Vision Transformers (ViTs) [11, 28, 44, 27] are now ubiquitous in computer vision. They decompose an image into a set of patches which are fed as tokens to a transformer model [43] of two main components: a spatial attention module, which reweighs each token according to its similarity to the other tokens extracted from the image, enabling information fusion across large spatial distances, and a channel mixer that combines the feature channels extracted from all patches using a multi-layer perceptron (MLP or FFN). A bottleneck of this model is the quadratic complexity of the attention mechanism on the number of patches. Numerous ViT variants have been proposed to address the problem, using improved attention mechanisms or hybrid architectures that replace attention or combine it with convolutions. Much less research has been devoted to the channel mixer. Most models simply adopt the two-layer MLP block of [43], where channels are first expanded by a specified expansion ratio and then compressed to the original dimension. This is somewhat surprising since the mixer is critical for good transformer performance. For example, it is known that 1) pure attention, without MLPs or residual connections, collapses doubly exponentially to a rank one matrix [10], 2) training fails to converge without residual connections or MLP [55], and 3) replacing MLPs with more attention blocks (both spatial and channel attention) of equivalent computational complexity lowers the transformer accuracy [8]. All these observations indicate that the channel mixer is an indispensable ViT component.
In this work, we uantify how much the channel mixer module contributes to ViT performance and investigate how to improve the trade-off between complexity and accuracy of the ViT model. We show that enhanced design of the channel mixer can lead to significant improvements in transformer performance by introducing a novel Scalable CHannEl MixEr (SCHEME) that enables the design of models with larger expansion ratios. SCHEME is a generic channel mixer that can be plugged into existing ViT variants to obtain effective scaled-down or scaled-up model versions. We replace the channel mixer of a state of the art architecture (SOTA) for low-complexity transformers, the MetaFormer-PPAA-S12 [55], with SCHEME to obtain a new family of SCHEMEformer models with improved accuracy/complexity trade-off. This is illustrated in Figure 1, where the SCHEMEformer family is demonstrated to establish new Pareto frontiers for accuracy vs FLOPS, accuracy vs model size, and accuracy vs throughput, showing that SCHEME allows fine control over all these variables, while guaranteeing SOTA performance. These properties are shown to hold for image classification, object detection and semantic segmentation tasks, as well as for different architectures, such as T2T-ViT [57], CoAtNet [5], Swin Transformer [27], CSWin [9] and DaViT [8].
To develop SCHEME, we start by studying the impact of the mixer channel expansion. Transformer performance is shown to increase with expansion ration until it saturates for an expansion ratio beyond 8. However, because mixer (MLP) complexity also increases with the dimension of the intermediate representation, naive channel scaling with expansion ratios larger than 4 causes an explosion of parameters and computation, leading to models of large complexity and prone to overfit. To achieve a better trade-off between dimensionality and computation, we leverage recent findings about the increased hardware efficiency of mixers with block diagonal structure [1, 6]. These group the input and output feature vectors of a layer into disjoint subsets and perform matrix multiplications only within each group, as illustrated in Figure 2. We denote the resulting MLP block as Block Diagonal MLP (BD-MLP). Despite the lack of feature mixing across groups, we find that a transformer model equipped with the BD-MLP and a larger expansion ratio (to match both the parameter count and computation) achieves comparable or slightly higher accuracy than a baseline with dense MLP. This suggests that the lower accuracy of the block diagonal operations is offset by the gains of larger expansion ratios. Further analyzing the features learned by the different groups of the BD-MLP, we observe that they form feature clusters of similar ability to discriminate the target classes. To learn better feature clusters, we seek a mechanism capable of restoring inter-group feature communication during training without increasing parameters. For this, we propose a channel attention branch that reweighs the input features of the BD-MLP according to their covariance matrix, as illustrated in Figure 2. This attention mechanism is denoted as the channel covariance attention (CCA) block. The re-weighted features are then fused with the BD-MLP output by means of a weighted residual addition with learned weights $(\alpha,1-\alpha)$.
As shown in Figure 2, SCHEME combines the sparse block diagonal structure of the BD-MLP, and the parameter-free CCA attention module, to implement a channel mixer extremely efficient in terms of parameters. Ablations of the evolution of the fusion weight $1-\alpha$ learned in the CCA branch (see Fig. 2) over training show that it gradually decays to zero during training. This happens consistently across all layers of the model and across model architectures. Hence, while the CCA is important for the formation of good feature clusters during training, it can be removed at inference without any loss, as illustrated in Table 3 and Figure 2. As a result, the model accuracy improves over simply using the BD-MLP mixer, but inference complexity does not. This leads to an extremely efficient inference setup, both in terms of parameters and FLOPs. Overall, the paper makes the following contributions,
@@
  • a study of the channel mixer of ViT MLPs, showing that dense feature mixing can be replaced by sparse feature mixing of higher internal feature dimensionality for improved accuracy, without increased complexity.
@@
  • the SCHEME module, which combines 1) a BD-MLP to enable internal feature representations of larger dimensionality than previous MLP blocks, and 2) CCA to enable the learning of these representations without cost at inference.
  • various models that combine SCHEME with previous transformer architectures to achieve SOTA trade-offs between accuracy and model size, FLOPS, or throughput, such as the SCHEMEformer of Figure 1. This is shown particularly effective for the design of fast transformers with small model size, of interest for edge devices, robotics, and low-power applications.
@@
  • Experiments on image classification, object detection, and semantic segmentation, showing consistent gains in accuracy for fixed computation and size.
@@ 2 Related Work
Vision Transformers: Vision transformers advanced the SOTA in several vision tasks since [28, 11]. This has quadratic complexity in the number of patches and is quite intensive. Most follow up work [17, 35, 56, 40, 36, 48, 37, 57, 58, 42, 38, 39, 44, 27, 42, 29, 61, 3, 18, 15] improved the spatial attention mechanism of ViT. A more extensive review of ViTs is given in [26].
@@
Role of attention: Recently, several works have shown that spatial attention is not the critical ViT feature. Some works improved performance by relying on hybrid architectures, which augment or replace ViT layers with convolutions [47, 41]. Other works have questioned the need for spatial attention altogether. [25] showed that spatial attention is not critical for vision transformers by proposing a spatial-gating MLP of comparable performance to ViT. Similarly, Metaformer [55] argued that the fundamental trait of ViT is the mixing of information across patches, showing that competitive results can be obtained by simply replacing attention with pooling or identity operations. DaViT [8] enhances performance by cascading spatial and channel attention blocks but increases transformer complexity, adding parameters and redundant channel mixing operations.
Role of MLP: Despite all this work on ViT architectures, little emphasis has been devoted to the channel mixer module (MLP) that follows attention. This is surprising because the mixer dominates both the parameter count and complexity (FLOPs) of the standard transformer block. [24, 56] modify the MLP block to mimic the inverted residual block of the MobileNetV2 [31], by adding a depthwise convolution. This improved performance but increases parameter and computation costs. Switch Transformer [14] replaces the FFN with a sparse mixture of experts [34] that dynamically routes the input tokens. This design allows scaling models to large sizes using higher number of experts but is not effective for ViTs. XCiT [12] employs a cross-covariance attention (XCA) operator, which can be seen as a "transposed" version of self-attention that operates across feature channels. However, it still employs the standard feedforward network after the XCA operation. Some recent works [52, 13] focus on MLP modules either by increasing the non-linearity with an arbitrary GeLU function or including modules such as SwiGLU, LayerNorm and RoPE which increases the complexity and latency of the model. In this work, we augment the FFN or MLP with a modified Cross-Covariance Channel Attention (CCA) operation during training to facilitate feature mixing across different channel groups that is applicable to various ViTs including XCiT to improve the computation-accuracy trade-off.
@@ 3 The SCHEME module
Figure 2 depicts the proposed SCHEME module for feature mixing in ViTs. As shown on the left, the standard channel mixer consists of two MLP layers, which expand the dimensionality of the input features and then reduce it to the original size. Let $\mathbf{x}\in\mathbb{R}^{d\times N}$ be the matrix containing the $N$ $d$-dimensional input feature vectors extracted from $N$ image patches. The mixer computes an intermediate representation $\mathbf{z}\in\mathbb{R}^{Ed\times N}$ and an output representation $\mathbf{y}\in\mathbb{R}^{d\times N}$ according to
@@
where $\mathbf{W}_{1}\in\mathbb{R}^{Ed\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times Ed},\mathbf{b}_{1}\in\mathbb{R}^{Ed},\mathbf{b}_{2}\in\mathbb{R}^{d},$ ${\bf 1}_{N}$ is the $N$-dimensional vector containing ones as all entries, $\sigma(.)$ is the activation function (typically GELU), and $E$ is an expansion factor, typically 4.
Table 1: MetaFormer-S12 [55] ImageNet-1K validation accuracy vs MLP expansion ratio ($E$). {*}: results by author code.
Ratio FLOPs Top-1 Acc
Model ($E$) (G) (%)
Metaformer 8 4.14 81.6 (+0.6)
Metaformer 6 3.35 81.8 (+0.8)
Metaformer [55] 4 2.55 81.0 (+0.0)
Metaformer 2 1.77 78.9 (-2.1)
Metaformer 1 1.33 76.0 (-5.0)
[…]
Table 2: SCHEME parameters (P) and FLOPs (F). 44-e8 (12-e8) downscales (upscale) the MetaFormer model.
SCHEME
44-e8 12-e8
S12 P (M) 12 21
S12 F (G) 1.8 3.3
S24 P (M) 21 40
S24 F (G) 3.3 6.5
S36 P (M) 55 59
S36 F (G) 8.0 9.6
Table 3: Impact of removing CCA branch during inference.
CCA Params FLOPs Top-1 Acc
SCHEME Used  (M) (G) (%)
44-e8-S12 11.83 2.16 79.74
44-e8-S12 11.83 1.77 79.72
12-e8-S24 40.0 7.3 82.80
12-e8-S24 40.0 6.5 82.76
CoatNet-44-e8 17.80 3.83 80.70
CoatNet-44-e8 17.80 3.42 80.68
Swin-12-e8-T 36.93 7.00 81.69
Swin-12-e8-T 36.93 5.89 81.69
Table 3 details the impact of the mixer on overall transformer performance by evaluating the role of the expansion factor $E$ on the performance of Metaformer-PPAA-S12 [55] (Pooling, Pooling, Attention, Attention) architecture on ImageNet-1K. Classification accuracy increases from 76.0 to $81.8\%$, as $E$ ranges from $1$ to $6$, decreasing for $E=8$, which suggests overfitting. The table also shows that these gains are not trivial. The S18 model, which has no expansion ($E=1$) but more transformer layers and complexity comparable to that of the S12 model with $E=4$, has an accuracy $2.7$ points lower than the latter. In summary, for a given computation budget, it is beneficial to trade off transformer depth for dimensionality expansion in the channel mixer. This shows that this expansion is a critical component of the transformer architecture. On the other hand, naively scaling $E$ beyond 6 severely increases parameters and computation, leading to models that over-fit and are impractical for many applications.
@@ 3.1 Scalable Channel Mixer (SCHEME)
Block Diagonal MLP (BD-MLP): Block diagonal matrices have been previously used to efficiently approximate dense matrices [1, 6]. In CNNs, group channel operations are frequently used to design lightweight mobile models with improved accuracy-computation trade-off [22, 2, 31]. This consists of splitting the feature vectors of (1)-(2) into disjoint groups, e.g. $\mathbf{x}$ into a set of $g_{1}$ disjoint features $\{\mathbf{x}_{k}\}_{k=1}^{g_{1}}$ where $\mathbf{x}_{k}\in\mathbb{R}^{d/g_{1}\times N}$, and $\mathbf{y}$ into a set $\{\mathbf{y}_{k}\}_{k=1}^{g_{2}}$ where $\mathbf{y}_{k}\in\mathbb{R}^{Ed/g_{2}\times N}$. As illustrated in Figure 2, the MLPs of (1)-(2) are then implemented independently for each group, according to
@@
$\mathbf{y}_{k}=\mathbf{W}_{2,k}\mathbf{z}_{k}+\mathbf{b}_{2,k}{\bf 1}^{T}_{N}$ (4)
where $\mathbf{W}_{1,k}\in\mathbb{R}^{Ed/g_{1}\times N/g_{1}},\mathbf{W}_{2,k}\in\mathbb{R}^{d/g_{2}\times Ed/g_{2}},\mathbf{b}_{1}\in\mathbb{R}^{Ed/g_{1}},\mathbf{b}_{2}\in\mathbb{R}^{d/g_{2}}$ and $\mathbf{z}$ is decomposed into a set $\{\mathbf{z}_{k}\}_{k=1}^{G}$ where $\mathbf{z}_{k}\in\mathbb{R}^{Ed/G\times N}$, with $G=g_{1}$ in (3) and $G=g_{2}$ in (4). Since the complexity of (3) is $g_{1}^{2}$ times smaller than that of (1) and there are $g_{1}$ groups, the complexity of the first MLP is $1/g_{1}$ times that of standard MLP. Similarly, the complexity of the second MLP is $1/g_{2}$ times that of the standard MLP. Hence, a transformer equipped with the BD-MLP and expansion factor $\frac{2g_{1}g_{2}}{g_{1}+g_{2}}E$ has identical complexity to a standard transformer of factor $E$. For example, when $g_{1}=g_{2}=g$ this allows growing the expansion factor by a factor of $g$ without computational increase.
@@
where the softmax operation is applied across the matrix rows and $\tau$ is a smoothing factor empirically set to 1 in our experiments. The output of the channel mixer block is the weighted summation of the BD-MLP and CCA branches
$\begin{split}\textbf{y}_{out}=\alpha\textbf{y}+(1-\alpha)CCA(\textbf{x}),\end{split}$ (6)
@@
CCA as a Regularizer: The introduction of a parameter free attention branch and a learnable weight $\alpha$ allows the model to form better feature clusters during training and gradually decay the contribution from CCA branch once the feature clusters are formed. This can be seen in Figure 4, which plots the value of the learned mixing weight $1-\alpha$ as a function of training epochs on ImageNet-1K, for all transformer layers. These plots are typical of the behavior we observed with all transformer backbones and architectures we considered. Clearly, $1-\alpha$ starts with high to intermediate values, indicating that information flows through both branches of the mixer, but decays to $1-\alpha\approx 0$ as training converges. Hence, as shown in Table 3, there is no degradation if the CCA branch is removed during inference. This eliminates a substantial amount of computation during inference, leading to the the training and inference setup of Figure 2, where CCA is not used at inference.
We explain this behavior by conjecturing that the downside of the computational efficiency of the BD-MLP is a more difficult learning problem, due to the independent processing of channel groups. This creates symmetries in the cost function, e.g. the order of the feature groups is not important, and requires a feature clustering operation that is likely to produce more local minima. The CCA branch helps to smooth out this cost function during training, while the feature groups are not established, by allowing inter-group communication. However, once the right feature groupings are found, CCA is no longer needed, and a simple BD-MLP mixer has no loss of performance over the standard MLP. Note that the operation of (5) is basically a projection of $\mathbf{x}$ into canonical subspaces of features that are correlated in the input image. This is likely to be informative to guide the group formation, but less useful when the features are already clustered. While this hypothesis is not trivial to test, since (5) varies from example to example, we confirmed that using CCA during training enhances class separability, which likely reduces overfitting for large expansion ratios. See section 4.2 for more details.
Computational Complexity: The complexity of the BD-MLP block is controlled by the group numbers $g_{1}$, $g_{2}$ and the expansion factor $E$, with a total cost $\mathcal{O}(Ed^{2}/g_{1}+Ed^{2}/g_{2})$, where $d$ is the channel dimension. The computational cost of CCA is $\mathcal{O}(Nd^{2})$ where $N$ is the number of tokens. Since CCA is not used during inference, it only adds to the computations during training. SCHEME provides a systematic framework that balances transformer width and depth by adjusting block size and expansion hyperparameters.